OpenStudy (anonymous):

The sum of three numbers is 7. One of the numbers is one more than the sum of the other two numbers. It is also four times the difference between the two numbers.

OpenStudy (anonymous):

Let the number be a , b, and c. a + b + c = 7 c - 1 = a + b c = 4( b -a ) from first and second equation , (a + b) + c = 7 (c - 1) + c = 7 2c - 1 = 7 => c = (7+1)/2 =4 so a + b =7- 4 = 3 and 4(b-a) = 4 divide second by 4 and add both the equations.. a + b = 3 b - a = 1 => 2b = 4 => b = 2 => a = 1 Hence your numbers are 1 , 2 , 4 .

OpenStudy (anonymous):

you understand the answer?

OpenStudy (anonymous):

yes thank you